### v1

# Reciprocal Reward Influence Encourages Cooperation From Self-Interested Agents

Abstract
Emergent cooperation among self-interested individuals is a widespread phenomenon in the natural world, but remains elusive in interactions between artificially intelligent agents. Instead, naïve reinforcement learning algorithms typically converge to Pareto-dominated outcomes in even the simplest of social dilemmas. An emerging class of opponent-shaping methods have demonstrated the ability to reach prosocial outcomes by influencing the learning of other agents. However, they rely on higher-order derivatives through the predicted learning step of other agents or learning meta-game dynamics, which in turn rely on stringent assumptions over opponent learning rules or exponential sample complexity, respectively. To provide a learning rule-agnostic and sample-efficient alternative, we introduce Reciprocators, reinforcement learning agents which are intrinsically motivated to reciprocate the influence of an opponent’s actions on their returns. This approach effectively seeks to modify other agents’ $Q$-values by increasing their return following beneficial actions (with respect to the Reciprocator) and decreasing it after detrimental actions, guiding them towards mutually beneficial actions without attempting to directly shape policy updates. We show that Reciprocators can be used to promote cooperation in a variety of temporally extended social dilemmas during simultaneous learning.
1 Introduction
Many species exhibit cooperative behaviors of remarkable variety and complexity. Even among prosocial animals, however, humans are especially notable for their ability to cooperate with unrelated individuals and maintain that cooperation even in highly adversarial environments (Melis & Semmann, 2010). These qualities are often credited with the development of human technology, culture, and advanced cognition (Burkart et al., 2014). As artificially intelligent (AI) agents are integrated into human society, it is increasingly important to ensure that they share these same prosocial qualities so that interactions between AI agents, as well as between AI agents and humans, may converge to mutually beneficial outcomes.
However, state-of-the-art reinforcement learning (RL) methods are typically designed for the single agent setting and are ill-suited for the nonstationarities introduced by multiple agents learning simultaneously. Treating other agents as fixed elements of the environment, referred to as the “naïve” learning (NL) approach, can destabilize training and produce globally suboptimal outcomes. This is particularly evident in a class of games known as sequential social dilemmas (SSDs), which contain tensions between collective and individual return (Leibo et al., 2017). SSDs present a particularly challenging problem for multi-agent RL (MARL) because of this mixed motivational structure, which precludes the use of common centralized training algorithms designed for fully-cooperative settings (Kraemer & Banerjee, 2016; Sunehag et al., 2017; Gupta et al., 2017; Rashid et al., 2018). Previous work has shown that NL agents optimizing only for their individual returns converge to non-cooperative, Pareto-dominated outcomes in even the simplest of SSDs, including mutual confession in the iterated Prisoners’ Dilemma (IPD) and an expected reward of 0 in Coins (Foerster et al., 2018).
In this work, we use an intrinsic reward to encourage agents called Reciprocators to reciprocate the influence that others exert on their returns, with the goal of inducing mutually beneficial behavior from other agents. We first define a notion of influence that quantifies the effect that one agent’s action has on another’s expected return. Then, given a pair of agents, a Reciprocator $rc$ and another agent $i$, we track the cumulative influence of agent $i$’s sequence of actions on $rc$’s expected return, which we call the influence balance that agent $i$ owes $rc$. At any time step, the Reciprocator receives an intrinsic reciprocal reward proportional to the product of its influence balance against agent $i$ and the influence of its selected action on agent $i$’s return. This incentivizes the Reciprocator to take reciprocating actions whose influence matches the sign of its current influence balance. For example, if agent $i$ has a net positive influence balance with $rc$, then $rc$’s intrinsic reciprocal reward will be positive when taking actions that positively affect agent $i$’s return.
The contributions of this work are as follows: (1) We formulate a novel intrinsic reward that encourages an agent to incentivize cooperation by influencing others’ expected returns without modifying the reward structure of the environment, requiring higher-order derivatives, or learning meta-game dynamics. (2) We show that agents trained with this reward achieve state-of-the-art cooperative outcomes in sequential social dilemmas without needing training schedules to control opponent selection. (3) Agents equipped with our intrinsic reward are able to induce cooperative behavior from purely self-interested naïve learners and several higher-order baselines.
2 Related Work
In order to improve convergence towards Pareto-optimal solutions among independently learning agents, previous work has modified agents’ reward structures to explicitly consider either group or per-capita return in order to promote cooperative behavior, e.g., via inequity aversion (Hughes et al., 2018), “empathic” harm reduction (Bussmann et al., 2019), or “altruistic” gradient adjustments (Li et al., 2024). While agents that abide by such restrictions may seek to reduce their harmful externalities, they have no way of regulating the behavior of other, less magnanimous agents, including purely self-interested “exploiters” in the worst case (Agapiou et al., 2023). Their efficacy in SSDs, as well as in most other types of multi-agent systems, is therefore limited in the absence of strong assumptions over the other agents’ altruistic tendencies.
Another avenue towards a more robust form of cooperation relies on the social mechanisms of reward and punishment, which are believed to assist in stabilizing cooperative relationships by controlling free-riding, cheating, and other antisocial behaviors (Melis & Semmann, 2010). Incorporating these mechanisms into groups of agents, whether by providing rewards to incentivize good behavior (Yang et al., 2020), or doling out punishments to discourage bad behavior (Schmid et al., 2021; Yaman et al., 2023), has been shown to encourage cooperation among independent agents in a variety of sequential social dilemmas (SSDs). However, these methods require extensions of the original action space that allow agents to directly modify other agents’ rewards. On the other hand, we instead propose to quantify the influence of each naturally available action $a\in A$ in a given state $s$ on another agent’s expected return $R$, and use this influence as the medium for reciprocation. As we will show, such reciprocal reward and punishment can be seen as a form of “opponent shaping” that seeks to manipulate the $Q$-value of given state-action pair and consequently the likelihood of that action in future policies.
Considering future policies points to a key issue with the canonical reinforcement learning (RL) framework, in that it optimizes only for the expected return within a single episode of the environment. However, taking actions that seek to optimize the long-term behavior of the other agents in the environment will not receive any immediate feedback within an episode, and must wait for one, or possibly many, learning steps. Opponent shaping methods of this kind address this issue by differentiating across pairs of episodes (Yang et al., 2020, LIO), through opponents’ gradient updates (Foerster et al., 2018, LOLA), or repeated sequences of learning steps organized into “meta-episodes” (Lu et al., 2022, MFOS). While these methods have demonstrated convergence to cooperative behavior in simple SSDs, they are impractical or intractable in realistic multi-agent scenarios, requiring additional independent action channels for providing incentives, white-box access to the learning rules and gradients of other agents, or exponential sample complexity to learn the dynamics of meta-games, respectively.
In particular, LOLA and subsequent works such as Proximal Learning with Opponent-Learning Awareness (Zhao et al., 2022, POLA) and Consistent Learning with Opponent-Learning Awareness (Willi et al., 2022, COLA) differentiate through the opponent’s learning update but have only been demonstrated to work with vanilla gradient descent and fixed learning rates. However, modeling the learning rule of each opponent becomes increasingly implausible in the face of modern RL algorithms, which uses adaptive optimizers that set different per-weight learning rates (Kingma & Ba, 2017, Adam), randomized experience replay (Schaul et al., 2016), and various auxiliary terms such as policy divergence penalties (Schulman et al., 2017) and entropy exploration bonuses (Williams, 1992; Mnih et al., 2016). Indeed, LOLA with opponent modeling (LOLA-OM) showed significantly worse results than agents with direct access to their opponent’s learning rule. Another class of opponent-shaping algorithms such as MFOS uses meta-learning approaches, which must repeatedly test new meta-policies in full training runs known as “meta-episodes” in order to learn how to exploit opponent learning dynamics. This requires the ability to freely perform rollouts in the environment against opponents whose policies can be reset to initialization at will, an unrealistic assumption in environments with partially adversarial motivations.
Our work is conceptually similar to Learning with Opponent Q-Learning Awareness (Aghajohari et al., 2024, LOQA), which also seeks to shape opponent policies by influencing the $Q$-values for different actions, under the assumption that opponents are Q-learners. However, LOQA still differentiates through a model of the opponent’s policy and optimizes according to the joint advantage function $A(s_{t},a_{1},a_{2})$ computed with respect to the state-value function $V(s_{t})$. On the other hand, we use a counterfactual baseline that marginalizes out only the opponent’s action to perform agent-specific credit assignment and use an intrinsic reward instead of gradients over opponent policies to encourage agents to influence the opponent’s $Q$-values in the correct direction. Most importantly, LOQA focuses on the problem of learning a general end-policy that performs well against a variety of other agents at evaluation and has therefore only demonstrated the ability to shape other LOQA opponents in a controlled self-play scenario, while our method focuses on shaping the policies of diverse agents over multiple episodes of simultaneous learning.
In the context of these higher-order opponent-shaping approaches, we position our intrinsic reciprocal reward as a form of immediate, within-episode feedback that encourages otherwise naïve learners to implicitly consider the long-term, cross-episode effects of their actions on the policy changes of other agents, demonstrating many of the properties of higher-order opponent-shaping methods while using only first-order reinforcement learning algorithms and standard rollout-based training procedures.
3 Preliminaries
We construct a series of sequential social dilemmas which can each be described as a stochastic game $G$, defined by a tuple $G=\langle S,A,P,r,n,\gamma\rangle$. In $G$, $n$ agents, indexed by $i\in\{1,\ldots,n\}$, observe the state of the environment $s\in S$ and simultaneously choose actions $a^{i}\in A$ to form a joint action $\boldsymbol{a}\in\boldsymbol{A}\equiv A^{n}$. The environment then undergoes a change according to the state transition function $P(s^{\prime}|s,\boldsymbol{a}):S\times\boldsymbol{A}\times S\rightarrow[0,1]$. Each agent receives an individual reward $r^{i}=r^{i}(s,\boldsymbol{a})$, and future rewards are discounted at each time step by a discount factor $\gamma\in[0,1]$. We consider the fully observable setting where agents observe the full state of the environment $s\in S$ at every time step, joint action $\boldsymbol{a}\in\boldsymbol{A}$, and rewards received by each agent. Each agent $i$ conditions a stochastic recurrent policy $\pi^{i}(a^{i}|\tau^{i})$ on its action-observation history, which is denoted as $\tau^{i}\in T\equiv(S\times\boldsymbol{A})^{*}$.
4 Reciprocal Reward Influence
4.1 1-Step Value Influence
Wang et al. (2020) proposed an intrinsic reward for exploration called Value of Interaction (VoI). VoI measures how much agent $i$’s actions affect agent $j$’s expected return by comparing the state-action value conditioned on the joint state and action $Q_{j}^{\boldsymbol{\pi}}\left({s},\boldsymbol{a}\right)$ against a counterfactual baseline $Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)$, where the $\boldsymbol{\pi}$ superscript indicates the dependence of this function on the parameters of all agents’ policies. This baseline marginalizes out the state and action of the influencing agent $i$ to compute a “default” expected return. This difference of $Q$-values can be decomposed into an immediate influence $r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})$ and a discounted future influence computed over changes in state transition probabilities [Equation 2]. Although the original VoI was formulated as a regularizer using an expectation over trajectories $\tau$, we modify it to compute the one-step Value Influence (VI) where $VoI=\mathbb{E}_{\tau}[VI]$, allowing us to compute the influence of individual actions
$VI_{i\mid j}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}\right)$ $=Q_{j}^{\boldsymbol{\pi}}\left({s},\boldsymbol{a}\right)-Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)$ (1)
$=r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})+\gamma\sum_{s^{\prime}}\left(1-\frac{p^{\boldsymbol{\pi}}\left(s^{\prime}\mid s,\boldsymbol{a}^{-i}\right)}{p\left(s^{\prime}\mid s,\boldsymbol{a}\right)}\right)V(s^{\prime}).$ (2)
This definition of influence relies on the notion of a counterfactual baseline to assign credit to a particular agent’s action while holding all other agents’ actions constant. This counterfactual baseline is so named because it estimates the counterfactual expected return if the agent’s true action is replaced with a “default” action, which is computed by marginalizing out the agent’s action to get the expected on-policy return. This baseline return is given by
$Q_{j\mid i}^{\boldsymbol{\pi}}\left(s,\boldsymbol{a}^{-i}\right)=\sum_{a_{i}}\pi^{i}(a_{i}|\tau)Q_{j}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i})).$ (3)
In practice, we approximate this by regressing towards the Q-value while masking out $a_{i}$ from the joint state-action input. The concept of marginalizing out an agent’s action as a baseline to quantify influence has also been used to define 1-step adversarial power (Li & Dennis, 2023), which estimates the maximum reduction in agent $i$’s expected reward that can be achieved by agent $j$’s action by taking the minimum over $a^{j}\in A$ in Equation 1. Similarly, Counterfactual Multi-Agent policy gradients (Foerster et al., 2017, COMA) marginalizes out a single agent’s action to assess the advantage (i.e., influence) of that agent’s selected action relative to the counterfactual on-policy return ceteris paribus, which is a special case of $VI_{i|j}$ where $i=j$ (Equation 4).
$A_{i}(s,\mathbf{a})=Q_{i}(s,\mathbf{a})-\sum_{a_{i}}\pi^{i}(a_{i}|\tau)Q_{i}^{\boldsymbol{\pi}}(s,(\boldsymbol{a^{-i}},a^{i})).$ (4)
4.2 Keeping Score with Influence Balances
The amount of influence that a Reciprocator is able to exert in a single time step is heavily environment-dependent, so that it may not always be possible to immediately reciprocate previous influences. To encourage reciprocation over extended timescales, we continuously accumulate a measure of net influence over sequences of actions. We draw inspiration from the “debit” mechanism $D_{i}$ used by approximate Markov tit-for-tat (Lerer & Peysakhovich, 2018, amTFT) to accumulate the advantage $A$ gained by an opponent $i$ compared to a cooperative counterfactual baseline, which they compute by recursively summing $A$ at every timestep.
Extending this idea to our framework, we sum agent $i$’s influence $VI^{\boldsymbol{\pi}}_{i|rc}$ on the Reciprocator’s expected return at each timestep rather than its own. However, using influence in only one direction as motivation for reciprocation can lead to continuous punishment or rewarding without a way to settle the score. To mitigate this, we also accumulate the net influence $VI^{\boldsymbol{\pi}}_{rc|i}$ in the opposite direction and subtract it from the influence balance at every timestep as a way to “pay off” the balance, limiting the degree to which reciprocation is rewarded. Formally, we define the influence balance $B_{rc|i}(t)$ maintained by a Reciprocator $rc$ with another agent $i$ at time $t$ as
$B_{rc|i}(t)=B_{rc|i}(t-1)+[VI^{\boldsymbol{\pi}}_{i|rc}(s_{t},\boldsymbol{a}_{t})-VI^{\boldsymbol{\pi}}_{rc|i}(s_{t},\boldsymbol{a}_{t})].$ (5)
The influence balance can be thought of as a score of net influence over time between agents, and can be used to motivate reciprocation in the correct direction, i.e., either positive reinforcement of net positive influence or positive punishment of net negative influence.
4.3 Intrinsic Reciprocal Reward
If agent $i$ takes a series of actions that improves the expected return of the Reciprocator over the learned baseline expectation, i.e., produces a net positive influence balance, then the Reciprocator should be motivated to reinforce this behavior by exerting a reciprocal positive influence on agent $i$’s expected return $R_{i}$, in order to encourage a higher likelihood of that behavior during policy updates. Similar logic applies for detrimental deviations, negative influence balance, and subsequent reciprocal punishment. We then define the intrinsic reciprocal reward $r^{R}_{rc|i}(t)$ received by $rc$ as
$r^{R}_{rc|i}(t)=B_{rc|i}(t-1)\cdot VI_{rc\mid i}^{\boldsymbol{\pi}}\left(s_{t},\boldsymbol{a}_{t}\right).$ (6)
Taking the product of existing influence balance and current action’s VI encourages the Reciprocator to take actions that reinforce agent $i$’s behavior in the correct direction by matching signs, and scales the reward by the magnitude of the outstanding influence balance and the reinforcing influence. The intrinsic reward is then added to the agent’s extrinsic reward to form the total reward used in training.
Algorithm 1 Training with Reciprocal Reward Influence vs. Agent $i$
  Initialize agent $rc$’s policy parameters $\theta_{\pi}$, VI target function parameters $\phi$, influence balance vector $B_{rc|i}=\boldsymbol{0}$, policy memory $\mathcal{M}=\emptyset$, and influence memory $\mathcal{H}=\emptyset$
  for each episode $e$ do
     Observe initial state $\boldsymbol{s_{0}}$
     for $t=1$ to $T$ do
        Choose action $a_{t}\sim\pi_{\theta}(s_{t})$
        Observe $\boldsymbol{a}_{t},\boldsymbol{r_{t}},\boldsymbol{s}_{t+1}$
        Store transition tuple $(s_{t},a_{t},r_{t},s_{t+1})$ in $\mathcal{M}$ and joint transition tuple $(\boldsymbol{s_{t}},\boldsymbol{a_{t}},\boldsymbol{r_{t}},\boldsymbol{s_{t+1}})$ in $\mathcal{H}$
     end for
     Update influence target function parameters $\phi$ with $\mathcal{H}$ every $k$ episodes
     Compute reciprocal rewards $r^{R}_{rc|i}(1),\ldots,r^{R}_{rc|i}(T)$ w.r.t. agent $i$ and sum with rewards in $\mathcal{M}$
     Compute advantage estimates $\hat{A}_{1},\ldots,\hat{A}_{T}$
     for $K$ epochs do
        Optimize the surrogate PPO-clip objective w.r.t. $\theta_{\pi}$
     end for
     Reset $\mathcal{M}=\emptyset$
  end for
4.4 Policy Optimization with Reciprocal Rewards
We use experience replay to periodically update target networks and iteratively update our counterfactual baseline estimates towards these target values (Mnih et al., 2015). This provides two key benefits: first, periodically updating these policy-dependent functions stabilizes training and allows us to approximately ignore their gradients, and therefore the gradient of the intrinsic reciprocal reward, with respect to the agents’ policy parameters (Wang et al., 2020). With this assumption, we are able to use to use standard policy gradient methods to train our agents to jointly optimize the combined extrinsic and intrinsic rewards.
Second, drawing samples from multiple previous episodes to the train the counterfactual baseline target functions makes the Reciprocator less susceptible to exploitation. If the counterfactual estimators were updated concurrently with the policy using only the most recent on-policy data, then the Reciprocator’s baseline would be immediately adjusted to its opponent’s new policy after each episode. Because assessment of influence hinges on these counterfactual baselines, updating them too frequently would allow adversaries to easily manipulate these estimates of on-policy returns.
5 Experiments
We conduct experiments using two commonly-used SSDs of varied complexity to demonstrate the shaping abilities of Reciprocators against other types of learning agents. For IPD, we consider memory-1 iterated games as in Foerster et al. (2018) and Lu et al. (2022), following the proof from Press & Dyson (2012) that longer-memory strategies provide no advantage over strategies conditioned on shorter memories. We use two methods to evaluate head-to-head performance in IPD: allowing agents to directly differentiate through the analytic, closed-form solution of the game as originally derived in Foerster et al. (2018), and more standard batched policy rollouts for a fixed episode length. We append “analytic” and “rollout” to the game names to denote the evaluation method used. For Coins, we augment the observation with the time remaining in the episode to prevent state aliasing and stabilize learning from experience replay (Pardo et al., 2022).
5.1 Sequential Social Dilemmas
Iterated Prisoners’ Dilemma (IPD): The iterated prisoners’ dilemma (IPD) is a temporally-extended version of the classical thought experiment, in which two prisoners are given a choice to either stay silent/cooperate (C) or confess/defect (D) with rewards given in Table 1(a). The Pareto-optimal strategy is for both agents to cooperate by maintaining their silence, but the only Nash equilibrium (in the non-iterated, single-shot case) is mutual defection.
Coins: Coins is an $n$-player game introduced by Lerer & Peysakhovich (2018) as a temporally-extended version of the IPD. In this game, $n$ players move around in a fixed-size grid and collect colored coins. Agents receive a reward of 1 for collecting a coin, and are punished if another agent collects their coin such that, if agents collect coins indiscriminately where $P(\text{collect own coin})\approx 1/n$, the net expected reward is 0. We show example rewards for $n=2$ in Figure 1(b).
C D
C -1, -1 -3, 0
D 0, -3 -2, -2
(a) Rewards for IPD.
(b) Rewards for Coins.
Figure 1: (a) The first number in each cell denotes the reward received by the agent taking the row action, and the second the reward received by the agent taking the column action, where C: cooperate (stay silent) and D: defect (confess). (b) Two agents (red and blue) are tasked with collecting randomly spawning coins. If an agent collects its own coin, it receives a reward of +1 (left). If an agent collects another’s coin, then it receives a reward of +1 but the other agent receives a punishment of -2.
5.2 Baselines
Although we have discussed an extensive body of work related to cooperation in SSDs, we select these baselines because they focus specifically on the problem of shaping other agents’ policies during simultaneous learning.
Naïve Learner (NL): As previously defined, NLs optimize their expected return with respect only to their own policy parameters $\theta$. In this work, we implement NLs using standard policy gradient methods (Sutton et al., 1999), which perform updates of the form $\theta_{t+1}=\theta_{t}+\alpha\nabla_{\theta}J(\pi_{\theta})|_{\theta_{t}}$, where $\alpha$ is the learning rate and $\nabla_{\theta}J(\pi_{\theta})|_{\theta_{t}}$ is the gradient of the objective function with respect to the policy parameters $\theta_{t}$ at step $t$.
Table 1: IPD-Analytic round robin results
RC NL LOLA M-MAML MFOS
RC -1.06 -1.03 -1.05 -1.05 -1.06
NL -1.06 -1.98 -1.52 -1.28 -1.88
LOLA -1.08 -1.30 -1.09 -1.04 -1.02
M-MAML -1.13 -1.25 -1.15 -1.17 -1.56
MFOS -0.98 -0.65 -1.02 -0.81 -1.01
Each entry is the average reward per episode achieved by the row agent against the column agent. We bold cooperative outcomes, which we define as both agents’ outcomes being better than -1.25, with a difference of no more than 0.10 between them. Standard error of the mean is less than 0.01 for all experiments. M-MAML results are averaged across 10 initial policies.
Figure 2: Representative run of a Reciprocator vs. an NL in IPD-Rollout. Reciprocal reward per step (left axis) and percentage of cooperation (staying silent) over the course of an episode (right axis).
Learning with Opponent-Learning Awareness (LOLA): LOLA uses either whitebox access to an opponent’s gradients and Hessians or an explicit model of their policy parameters, assuming they are NLs, and differentiates through their learning step using the update rule
$\theta_{t+1}^{i}$ $=\theta_{t}^{i}+\alpha^{i}\nabla_{\theta_{t}^{i}}J^{i}\left(\theta_{t}^{i},\theta_{t}^{-i}+\Delta\theta_{t}^{-i}\right)$
$\Delta\theta_{t}^{-i}$ $=\alpha^{-i}\nabla_{\theta_{t}^{i}}J^{-i}\left(\theta_{t}^{i},\theta_{t}^{-i}\right).$
Multiagent Model-Agnostic Meta-Learning (M-MAML): M-MAML (Lu et al., 2022) learns initial parameters and then meta-learns over both its own and its opponent’s policy updates, conceptually similar to Meta-Multiagent Policy Gradient (Kim et al., 2021, Meta-MAPG) but modified to differentiate directly through the analytic form of the return in matrix games.
Model-Free Opponent Shaping (MFOS): As briefly discussed in the introduction, MFOS meta-learns over multiple episodes of policy updates in order to accomplish long-horizon opponent shaping. In Lu et al. (2022), MFOS is implemented with inner and outer policies, where the outer policy either directly outputs an inner policy to play in each episode or a conditioning vector which is elementwise multiplied with an inner policy vector, as done in IPD-Analytic and Coins, respectively.
5.3 Implementation Details
In IPD-Analytic, we differentiate directly through the analytic solution to the matrix game. For rollout-based experiments, we implement all policy gradient-based agents using actor-critic architectures trained with proximal policy optimization and a clipped surrogate objective (Schulman et al., 2017, PPO-Clip). Target networks to estimate the $Q$-values in Equation 1 are updated every $k$ episodes using uniformly sampled experience from a replay buffer (Lin, 1992). Additional hyperparameter values and network architecture details can be found in Appendix A.1. Shaded regions in plots depict one standard error of the mean.
6 Results
Here we summarize the results of our experiments, which include round-robin evaluations in IPD and demonstrations of Reciprocators’ ability to reach cooperation and shape NLs in Coins. Following work by Yu et al. (2022) showing that Meta-MAPG and LOLA-DICE agents fail to achieve significant results in Coins, even with a simplified shared reward, we do not use them as baselines in Coins.
6.1 IPD
IPD-Analytic: In the analytic form of IPD, we show that Reciprocators are able to reach cooperative equilibria with all other baselines despite being only a first-order method using vanilla gradient descent [Table 1]. Although the Reciprocator is extorted by MFOS to small extent (-1.06 vs. -0.98, respectively), we emphasize that MFOS relies on extensive liberties such as the ability to observe thousands of parallel training runs and roll out new meta-policies against opponents that are repeatedly reset back to randomly initialized policies. Additionally, as explained in Section 4, stochastic policy rollouts allow the Reciprocator to influence opponent returns differently for different action sequences and provide a stronger learning signal, motivating additional experiments using sampled rollouts.
IPD-Rollout: In IPD-Rollout, we show that Reciprocator is able to consistently shape an NL in IPD over the course of a single learning trajectory with limited rollout samples. We observe interesting learning dynamics wherein the Reciprocator’s policy repeatedly oscillates between lower and higher levels of cooperation. We show that these oscillations are driven by the opposing intrinsic and extrinsic rewards, and observe that the derivative of the rate of cooperation corresponds to the reciprocal reward [Figure 2]. Mechanistically, we see that the Reciprocator initially learns a cooperative tit-for-tat strategy guided by a strong reciprocal reward. However, as the NL’s policy becomes more deterministic and $a^{i}$ becomes predictable, the $VI^{\boldsymbol{\pi}}_{i|rc}$ component of the influence balance decreases. As the extrinsic reward begins to dominate, the Reciprocator essentially reverts into an NL, leading to exploitative defection. However, this exploitation of the NL agent $i$ causes the $VI^{\boldsymbol{\pi}}_{rc|i}$ term to become negative, leading to a negative reciprocal reward. This, combined with an increase in the NL’s frequency of defection (leading to an increase in $VI^{\boldsymbol{\pi}}_{i|rc}$), produces a reversal back to cooperative behavior that is then reinforced by positive reciprocal rewards.
Figure 3: Shaping an NL in Coins. Percentage of own coins collected by NL during training when facing each opponent (left) and coin counts across 32 steps for Reciprocator vs. NL (right).
6.2 Coins
In this temporally-extended social dilemma, we see that the Reciprocator is able to shape NL-PPO into picking up more of its own coins at rates equal to MFOS. This allows both agents to achieve a positive reward, while needing only a fraction of the samples to converge to this behavior (32,000 steps for Reciprocator vs. 512,000 steps for MFOS) [Figure 3]. We show that this change is driven by changes in coin preference rather than in total collection, with both agents collecting at near-optimal pace. This is in contrast to MFOS [Appendix B.2], which shapes the NL into collecting fewer coins.
When two Reciprocators are pitted against each other, we see that they are quickly able to learn a cooperative strategy of collecting their own coins [Figure 4], resulting in an average reward of $\sim 8$ per 32 steps without needing self-play. This significantly outperforms MFOS and the reported performance of LOLA with opponent modeling (LOLA-OM), which achieved an average reward of only $\sim 2$ per 32 steps according to Foerster et al. (2018). One concern is that the intrinsic reciprocal reward overpowers the extrinsic reward, effectively changing the motivational structure of the game into one that is fully-cooperative. To address this, we additionally recorded the cumulative reciprocal reward received in each episode over the course of training [Appendix B.1] to show that the reciprocal reward quickly decreases to a negliglible fraction of the total reward.
Figure 4: Head-to-head results in symmetric Coins (two agents of the same kind). Total reward per episode (left), percentage of own coins collected (right). Reciprocators and Naive Learner results are plotted on a scale of single episodes (bottom x-axis) whereas MFOS results are plotted on a scale of meta-episodes, where one meta-episode is equivalent to 16 episodes (top x-axis).
Together, these results demonstrate that Reciprocators are able to robustly shape the behavior of other agents towards prosocial equilibria during simultaneous learning, achieving state-of-the-art results with fewer assumptions and limitations than existing methods. Apart from opponent-shaping properties, we also show that the intrinsic reciprocal reward discourages Reciprocators from exploiting others, showing promise for the development of a more cooperative multi-agent learning framework.
7 Limitations and Future Directions
Our current implementation of Reciprocators using naïve RL algorithms remains limited in that it seeks to maximize its compound return within single episodes rather than across multiple episodes. Using handpicked weights to balance intrinsic and extrinsic rewards opens the door for a suboptimal tradeoff between long-term opponent shaping and short-term return maximization. This is partially addressed by the counterfactual target baseline updates, which allow the Reciprocator to lower the magnitude of the reciprocal reward in response to opponent policies that remain stationary over multiple episodes [Figure 2]. Future work will focus on methods to evaluate reciprocation efficacy across multiple episodes and dynamically tune the balance between reciprocal and extrinsic rewards.
8 Conclusion and Broader Impacts
Emerging interest in cooperative AI has been led by approaches that endow agents with higher-order shaping capabilities. Although these agents exhibit cooperative behaviors when pitted against equivalent opponents, they readily manipulate and exploit agents with simpler learning rules or lower computational capabilities. This is a fundamentally undesirable outcome in partially-adversarial interactions between learning agents, with the potential to exacerbate existing computational resource gaps beyond out-of-the-box performance.
We presented Reciprocators, naïve learning agents equipped with an intrinsic reward for multi-agent settings that seeks to influence the behavior of other, simultaneously learning agents towards mutually beneficial outcomes. We showed that Reciprocators are able to shape the behavior of other agents in the environment over the course of training, and can produce prosocial behaviors in a variety of sequential social dilemmas. To the best of our knowledge, Reciprocators comprise the first class of learning algorithms to achieve cooperation during simultaneous learning between two independent agents without needing meta-learning methods, knowledge of other agents’ learning algorithms, or training schedules such as self-play or tracing procedures which control opponent selection and rely on the symmetric player roles. We believe that these results show a promising avenue forward for inducing cooperative outcomes from a diverse array of learning algorithms.
Appendix A Appendix / supplemental material
All experiments were run on an Nvidia 3070 GPU with 8 GB of VRAM. For IPD-Analytic and Coins experiments involving MFOS agents, we adapted the original code from Lu et al. (2022).
A.1 Hyperparameter Details
Each experiment was run 8 times, with a batch size of 8192 for IPD-Analytic and 2048 for IPD-Rollout and Coins.
Hyperparameter IPD-Rollout Coins
Number of Linear Layers 2 2
Size of Linear Layers 2 16
Number of GRUs - 1
Size of GRUs - 16
Episode Length 32 32
Adam Learning Rate 0.005 0.005
PPO Epochs Per Episode $K$ 10 40
PPO-Clip $\epsilon$ 0.1 0.15
Discount Factor $\lambda$ 0.96 0.99
Entropy Coefficient 0.02 0.01
Table 2: General PPO parameters.
With the exception of a linear layer size of 32 for Coins, network parameters to estimate the various target functions to compute the VI are the same as those listed for the PPO network in Table 2. Target estimates for IPD policies were computed by averaging across observed choices for each of the five possible states (start, CC, CD, DC, DD). Replay buffer sizes are in units of episodes, where each episode consists of a batch of 32 steps. Then, for example, a replay buffer size of 4 for Coins corresponds to 32 steps $\times$ 4 episodes $\times$ 2048 batch size = 262,144 steps of experience stored in the replay buffer.
Hyperparameter IPD-Analytic IPD-Rollout Coins
Replay Buffer Size (in episodes) 5 1 4
Target Function Update Period 10 3 1
Target Function Epochs Per Episode - - 20
Adam Learning Rate - - 0.01
Reciprocal Reward Weight 5.0 5.0 1.0
Table 3: Reciprocator-specific parameters.
Again, network parameters for the outer meta-policy and inner actor and critic networks are the same as the general PPO parameters above.
Hyperparameter IPD-Analytic IPD-Rollout Coins
Outer Episode Length 16 16 16
Inner Episode Length - 32 32
Table 4: MFOS-specific parameters.
Appendix B Additional Experiments
B.1 Reciprocal Reward in Coins
An example run of two Reciprocators in Coins showing the relative magnitude of intrinsic reciprocal reward and extrinsic environmental rewards over the course of training.
Figure 5: Mean extrinsic reward (left), mean intrinsic reward (right) for two Reciprocators in Coins.
B.2 Coin Counts vs. NL-PPO
We show the total coin counts over the course of training runs vs. an NL-PPO in order to show that changes in P(Own Coins) are driven by changes in coin color preference rather than changes in the total number of coins collected.
Figure 6: Percentage of own coins collected by NL-PPO (left) and total number of coins collected by NL-PPO (right) vs. each baseline in Coins.